Consumer A has a utility function u(x,y) = xA + yA and an endowment of (x,y) = (25,5). Consumer B has a
utility function u(x,y) = min{xB,yB} and an endowment (x,y) = (25,45).
a. Carefully sketch the Edgeworth Box and indicate where the endowment is.
b. What is A’s utility and B’s utility if they each simply consumer their endowments?
c. Next, add the indifference curve for A and B, through their endowments in your Edgeworth Box.
d. Find a Pareto Efficient allocation where A receives the same utility as she gets from her endowment.
Answer #1

Solution P.Nol Total x = 25+ 05 - 50 Total y = 45+5 = 50 This Edgeworth box is square Sized Endowment point Es - (25,5) B UAP. 2 D) Allocation (x,y) for A = (15,15) U.Az 30 for.B, (x,y) = (30, 30) min (30,30) U-Ba - 30 No, v better off , welfare of